OpenStudy (radar):

\[2\pi \times radians=360^{0}\]

OpenStudy (radar):

Say you wanted to express 270 degrees 270/360 =x/2 pi radians 270(2 pi radians)=360 x x=(270*2 pi radians)/360 x=1.5 pi radians

OpenStudy (anonymous):

I'm sorry. I still dont understand :/

OpenStudy (radar):

It is a ratio. You memorize a relation between degrees and some value of I use 2 pi radian = 360 degrees (full circle) you might prefer pi=180 degrees. then you use the concept of ratios. Comprehende?

OpenStudy (radar):

Play with it, you will soon "get it"